OpenStudy (onthecly2):

Convert the equation from Polar to Rectangular form: 5= 1/ 4cos(theta)-3sin(theta)

OpenStudy (holsteremission):

Is it supposed to be \[5=\frac{1}{4\cos\theta}-3\sin\theta\quad\text{or}\quad5=\frac{1}{4\cos\theta-3\sin\theta}\]

OpenStudy (onthecly2):

typo.. the 5 is supposed to be an r. so: \[r=\frac{ 1 }{ 4\cos \theta-3\sin \theta }\]

Nnesha (nnesha):

\[x=r \cos \theta ~~~~~,~~~y=rsin \theta\]

OpenStudy (onthecly2):

a. \[y=\frac{ 4 }{ 3 }x-1\] b. y=4x+1 c. \[y=\frac{ 4 }{ 3 }x-\frac{ 1 }{ 3 }\] d. y=12x+3

myininaya (myininaya):

hint multiply both sides by 4 cos(theta)-3 sin(theta) then make the replacements @Nnesha mentioned

OpenStudy (sshayer):

divide both sides by r then put \[r \cos \theta=x,r \sin \theta=y \]

OpenStudy (marcelie):

hmm i got c. you must know these x= rcostheta y=rsintheta